OpenStudy (anonymous):

sqrt 20 - 4 sqrt 45 + sqrt 80

OpenStudy (anonymous):

\[\sqrt{20}-4\sqrt{45}+\sqrt{80}\]

OpenStudy (anonymous):

guess what will you do now

OpenStudy (anonymous):

answer is \[-6\sqrt{5}\]

OpenStudy (anonymous):

okay i am doing this @nadialabish94 \[\sqrt{20} - 4\sqrt{45} + \sqrt{80} = \sqrt{2*2*5} - 4\sqrt{3*3*5} + \sqrt{4*4*5}\] which is equal to \[2\sqrt{5} - 12\sqrt{5} + 4\sqrt{5}\] \[\sqrt{5}\left( 2-12+4 \right)\] \[-6 \sqrt{}\]

OpenStudy (anonymous):

its \[-6 \sqrt{5}\]

OpenStudy (anonymous):

So I did \[\sqrt{20}+\sqrt{80}\] first, now im doing \[-4\sqrt{45}+6\sqrt{5}\] I just dont know how to do \[-4\sqrt{45}+6\sqrt{5}\] ?